OpenStudy (anonymous):

what is the radius of a kitchen table with a circumfrence of 19.79 feet ?.......... use pi=3.14.............round ur answer to the nearest hundred

OpenStudy (anonymous):

\[Circumference = 2 \pi r\] \[19.79 = 2 \pi r\] \[r = (19.79)/(2 \pi)\] \[Area = \pi r^{2}\] \[Area = \pi (19.79/(2 \pi))^{2} = 31.18\]

OpenStudy (anonymous):

Woops stepped past the radius \[r = 3.15 ft\]